OpenStudy (dls):

\large lim_{x rightarrow 0} frac{2sinx-sin2x}{x^3}

OpenStudy (dls):

\[\large \lim_{x \rightarrow 0} \frac{2sinx-\sin2x}{x^3}\]

hartnn (hartnn):

write sin 2x as 2 sin x cos x

OpenStudy (dls):

already did

OpenStudy (dls):

\[\large \lim_{x \rightarrow 0} \frac{2sinx-2sinxcosx}{x^3}\]

hartnn (hartnn):

take sin x common

OpenStudy (dls):

\[\large \lim_{x \rightarrow 0} \frac{2sinx(1-cosx)}{x^3}\]

OpenStudy (dls):

already did

OpenStudy (dls):

stuck here

hartnn (hartnn):

write one of the x under sin x and use sinx/x =1

OpenStudy (dls):

2sinx or sinx?

OpenStudy (dls):

will i be left with?: 2(1-cosx)/x^2?

hartnn (hartnn):

yup.

OpenStudy (dls):

ill be left with 1 right?

hartnn (hartnn):

now multiply and divide by 1+cos x

OpenStudy (dls):

okay

OpenStudy (dls):

thenn?

hartnn (hartnn):

what is (1+cos x)(1-cos x)

OpenStudy (dls):

sin^2x

OpenStudy (dls):

2/2=1

hartnn (hartnn):

and u have x^2 in denominator, right ?

hartnn (hartnn):

right 2 is correct

hartnn (hartnn):

sorry 1

OpenStudy (dls):

:D yayy!!